# On Context Utilization in Summarization with Large Language Models

### v3

@@ Abstract
Large language models (LLMs) excel in abstractive summarization tasks, delivering fluent and pertinent summaries. Recent advancements have extended their capabilities to handle long-input contexts, exceeding 100k tokens. However, in question answering, language models exhibit uneven utilization of their input context. They tend to favor the initial and final segments, resulting in a U-shaped performance pattern concerning where the answer is located within the input. This bias raises concerns, particularly in summarization where crucial content may be dispersed throughout the source document(s). Besides, in summarization, mapping facts from the source to the summary is not trivial as salient content is usually re-phrased. In this paper, we conduct the first comprehensive study on context utilization and position bias in summarization. Our analysis encompasses 5 LLMs, 10 datasets, and 5 evaluation metrics. We introduce a new evaluation benchmark called MiddleSum on the which we benchmark two alternative inference methods to alleviate position bias: hierarchical summarization and incremental summarization [^1] We will share all code and data upon publication..
@@ 1 Introduction
Large language models (LLMs) have drastically transformed the landscape of NLP recently Brown et al. (2020). With instruction tuning Ouyang et al. (2022); Chung et al. (2022), LLMs made a major leap forward in conditional (prompted) content generation, and can generate satisfying outputs without the need to finetune on a task specifically. In abstractive summarization specifically, this approach has arguably opened a new paradigm; summaries generated by LLMs are highly fluent, grammatical and relevant (Goyal et al., 2022). Despite noticeably lower scores on automatic metrics such as ROUGE (Lin, 2004) or BERTScore (Zhang et al., 2019), summaries generated by LLMs are largely preferred by humans over summaries from state-of-the-art fine-tuned models like BRIO (Liu et al., 2022b, 2023c). In fact, on XSum, GPT-3.5 summaries are even on par with re-annotated human-written summaries, and much better than the dataset ground-truth, according to human evaluators (Zhang et al., 2023b). LLMs also show promising capability in evaluating summaries generated by other systems, including other LLMs (Fu et al., 2023; Luo et al., 2023; Shen et al., 2023a).
Despite this success, a few major technological bottlenecks remain with LLMs, including the maximum length of their context window. The standard context window length for open-source LLMs is 2k tokens (Brown et al., 2020; Scao et al., 2022; Penedo et al., 2023; Touvron et al., 2023a), which drastically limits their usefulness for long-input summarization (Shaham et al., 2022). Several techniques were proposed to extend the context window, including ALiBi (Press et al., 2021), LeX (Sun et al., 2022), position interpolation (Chen et al., 2023) and YaRN (Peng et al., 2023). While some of them claim up to 128k tokens processing capacity (Peng et al., 2023), it remains unclear how much such methods help on long-context summarization.
@@
In this work, we investigate in depth how LLMs use their context window in abstractive summarization. Unlike in question-answering, mapping facts in the output to a specific snippet in the source is not straightforward in abstractive summarization, due to the high-level of re-phrasing and compression. We conduct a large-scale study with 5 open-source LLMs, 10 datasets covering many aspects of summarization, and 5 highly diverse automatic metrics. Our contributions are threefold:
@@ 2 Experimental Setup
We experiment with several popular high-performing open-source instruction-tuned LLMs. Instruction-tuning datasets such as Flan (Wei et al., 2021) include some of the datasets we study: CNN/DM, XSum, SAMSum and Multi-News. We analyze the following models through HuggingFace transformers (Wolf et al., 2020):
  • Flan-UL2 is a 20B parameters encoder-decoder model pre-trained on 1T tokens. It is based on the UL2 20B model (Tay et al., 2022), with the addition of Flan-T5 (Chung et al., 2022) instruction fine-tuning. The context window is 2k.
  • Llama-2 (Touvron et al., 2023b) is a recently introduced powerful decoder-only model pre-trained on 2T tokens, ranging from 7B to 70B parameters, and with a 4k context window. We use the 7B and 13B models.
  • Xgen-7B (Nijkamp et al., 2023) is a 7B decoder-only model pre-trained on up to 1.5T tokens. It supports 8k tokens context window.
@@
We use the instruction-tuned (or chat) checkpoints for Llama-2, Xgen-7B and Mistral-7B. To run inference, we use the following prompt: Read the following text and summarize it: [text]. Summarize the above text in [n] sentences. Summary: where n is set to an average number of target sentences per dataset (see Appendix A). We infer all models in bfloat16 and sample summaries with top-k sampling (Fan et al., 2018) using $k=50$.
@@
Summarization evaluation is especially challenging in the LLM era, as most automatic metrics poorly correlate with human preferences (Goyal et al., 2022; Liu et al., 2022a). To get a broad picture of performance, we evaluate with metrics as diverse as possible. First, we consider reference-based metrics: ROUGE-2 (Lin, 2004), which measures bigram overlap, BERTScore (Zhang et al., 2019), which measures semantic similarity with BERT (Devlin et al., 2018) embeddings, and A3CU (Liu et al., 2023c), which extracts facts in the form of Atomic Content Units (ACUs) Liu et al. (2022a), and checks the presence of ACUs between prediction and reference. As reference-free metrics, we include SummaC (Laban et al., 2022), a leading factual consistency evaluation metric relying on entailment scores between pairs of source and summary sentences. We also leverage GPT-3.5 (gpt-3.5-turbo-0125) as a summarization evaluator, which is proven to be a strong natural language generation evaluator (Wang et al., 2023a; Shen et al., 2023a; Jain et al., 2023). We prompt the model with the source and generated summary (which fits in GPT-3.5’s 16k context window) and ask to output a score on a likert scale from 1 to 5. We refer to Appendix B for the full prompt template. For GPT-3.5 evaluation, we subsample 300 data points per dataset in order to reduce costs.
We report the performance of LLMs on all 10 datasets, alongside a comparison to SOTA, in Appendix C. FLan-UL2 dominates on standard-length datasets, but Llama-2-13B has the upper hand on the long-input ones. Performance itself is not our focus in this paper, but rather which position-related factors influence it. We discard Flan-UL2 on long-input datasets due to poor performance.
@@ 3.1 RQ1: Where in the source do LLMs take their information from?
We investigate summaries generated by LLMs, and map them to specific parts of the input. Unlike in question-answering or extractive summarization, mapping salient information from a summary to the source is not trivial in abstractive summarization.
@@
As seen in Fig. 1, all summarization datasets except XSum, and Reddit-TIFU show some lead bias: salient bigrams from the reference (orange curves) are more likely to be found at the beginning of the source. However, LLMs show a significantly stronger lead bias on all datasets: bigrams from LLMs summaries are much more likely to be found in the first 20% words of the source. It is especially striking on XSum (except for Flan-UL2), Reddit-TIFU, Arxiv, PubMed and GovReport. On XSum, Flan-UL2 closely matches the reference distribution, which we attribute to its better instruction tuning. Results in Appendix D confirm that bigram distribution for LLMs and references are statistically different (p-value of Kolmogorov-Smirnov test (Massey Jr, 1951) inferior to 0.001) in all cases except Flan-UL2 on XSum and SAMSum and Llama-2-7B on XSum. We conclude that LLMs focus on contents at the beginning of the source document(s).
@@ 3.2 RQ2: Where do LLMs look at within their context window?
In the previous experiment’s design, LLMs may not see the entire source in long-input summarization datasets, due to their limited context window, which is shorter than the source on the long-input datasets. We now focus on input information accessible to LLMs, and only consider salient information if it falls within the context window. Besides, since the same bigram may occur multiple times throughout the source, we adjust the methodology for saliency estimation. We align sentences in generated summaries to sentences in the context, following the procedure described in (Zhou et al., 2018) and also used in (Adams et al., 2023). Specifically, we greedily select source sentences until the ROUGE-1 F1 score between the set of selected source sentences and the summary stops increasing. The resulting set of source sentences forms a proxy of the visible salient input information being rephrased by the model when summarizing. We split each truncated source document into 10 bins of the same number of sentences, and map each aligned source sentence to its bin. Note that bins are not directly comparable across models, as context length varies across models.
As we can see in Fig. 2, sentences from the first 10% or last 10% of the input context are much more represented than others. A clear U-shape emerges on PubMed and SummScreenFD for all LLMs. This is intriguing knowing that Llama-2 LLMs and the other two LLMs have different context window lengths, and the last 10% of each context window may contain content of varying saliency. In other words, LLMs seem to be mostly re-phrasing information from the beginning or the end of their context window.
@@ 3.3 RQ3: Does LLMs performance depend on the position of salient information?
We examine performance changes with regard to this salient position. To do so, we compute the Spearman correlation coefficient between salient position and each evaluation metric in Table 2. A high absolute Spearman value means that summary quality (as measured by this metric) can change (and deteriorate) with the position of important information within the context. There are several takeaway findings from this Table. First, we notice that on standard-length datasets, reference-based evaluation metrics are negatively correlated to position of salient information. The correlation is only moderate, yet remarkably consistent across datasets (except Reddit-TIFU) and models. This is surprising, since such datasets fit entirely in context and are not affected by truncation. In contrast, reference-free metrics show either no significant or positive correlation to information position. For long-input datasets, the negative trend for reference-based metrics is confirmed. On these lengthy datasets, SummaC and GPT-3.5 tend to switch from positive to negative correlation, especially for Xgen-7B and Mistral-7B. We highlight that since GPT-3.5 itself is affected by the middle-curse from Liu et al. (2023a), it may not accurately evaluate summarization when salient content lays in the middle of the context. In light of these results, we conclude that LLMs’ summarization performance is sensitive to the position of salient information in the context window.
@@ 4.1 How is information in the middle treated?
In the first experiment, we vary the position of salient information throughout the input. We keep a single document (the abstract of the query paper on Multi-XScience, and the document with the highest BERTScore with the reference on Multi-News), and place it at position $j$ for $j\in\{1,\ldots,k\}$, using $k-1$ documents from a random data point for the other slots. The single relevant document is accompanied by a [RELEVANT] header, while the other documents have an [IRRELEVANT] header, and we prompt the LLM to only summarize the relevant document. For reference-free evaluation metrics, we use the single relevant document as source. In this experiment and the following, we also include a random baseline of shuffled inputs and model predictions. In Fig. 3, we see a noticeable drop in performance for all metrics when the salient document is not in the first or final position. Flan-UL2 seems to focus on the end of the context, Xgen-7B and Mistral-7B on the beginning, and Llama-2 models on both. Performance can fall quite below random range, especially for SummaC, confirming the worrying trend from Liu et al. (2023a).
A more fine-grained analysis with GPT-3.5 in Fig. 4 evaluating specific attributes (following the method in Adams et al. (2023)) reveals more details. Coherence and quality remain high and stable. In other words, the text outputs are of good quality. But for informativeness and attributability, the U-shape appears again: it shows that the LLMs are struggling to generate content sticking to the document inserted in the middle.
Dataset Model Input documents R-2 BS A3CU SummaC GPT-3.5
Multi-X Llama-2-7B All 7 4.64 82.83 5.88 54.56 4.17
First + last 4.62 82.82 5.78 47.00 4.50
First + 5 random + last 4.43 82.64 5.37 43.25 4.01
Random 1.22 81.14 3.08 30.30 1.14
@@
First + last 4.73 82.86 5.76 43.74 4.53
First + 5 random + last 4.61 82.80 5.72 46.42 4.31
Random 1.16 81.16 3.22 28.38 1.42
@@
First + last 5.01 82.73 5.86 49.08 4.45
First + 5 random + last 3.89 82.16 5.03 55.29 3.01
Random 1.25 80.74 3.21 27.77 1.29
@@
First + last 5.12 82.67 6.15 60.91 4.76
First + 5 random + last 4.45 82.40 5.35 58.19 4.06
Random 1.25 80.64 3.13 25.97 1.18
Multi-N Llama-2-7B All 5 10.76 85.04 19.06 60.09 4.00
First + last 9.50 84.43 15.88 54.52 3.80
First + 3 random + last 7.57 83.36 12.39 50.35 2.94
Random 1.52 80.22 3.11 33.34 1.92
@@
First + last 9.55 84.58 16.99 49.84 3.73
First + 3 random + last 8.27 83.79 14.84 50.09 3.18
Random 1.55 80.05 3.17 32.80 1.98
@@
First + last 7.82 83.27 14.18 51.59 3.60
First + 3 random + last 6.30 81.85 11.66 49.02 2.66
Random 1.52 78.99 3.48 37.62 1.86
@@
First + last 9.11 83.69 14.99 67.14 3.51
First + 3 random + last 6.59 81.66 12.50 52.14 2.45
Random 1.57 79.21 3.30 37.18 1.80
Table 3: Performance in multi-document summarization on Multi-XScience (7 documents) and Multi-News (5 documents) when infilling the middle of the context window with random documents. R-2 is ROUGE-2, BS refers to BERTScore.
@@
In the second experiment, we take the opposite approach, and put salient information at the beginning and the end, while the middle of the prompt is filled with noise. We keep the first and last documents, and fill the $k-2$ middle ones with random documents. We also run a baseline just using the first and last documents as input, expected to be close to the result with random documents in between. As displayed in Table 3, filling with random noise between the first and last document (which amounts to a prompt mostly irrelevant to the reference) leads to a moderate drop in performance. For instance, on Multi-XScience, with 5 random documents between the first and last, Llama-2-13B maintains 4.31 as GPT-3.5 score compared to 4.35 when using all 7 documents.

### v4

@@ Abstract
Large language models (LLMs) excel in abstractive summarization tasks, delivering fluent and pertinent summaries. Recent advancements have extended their capabilities to handle long-input contexts, exceeding 100k tokens. However, in question answering, language models exhibit uneven utilization of their input context. They tend to favor the initial and final segments, resulting in a U-shaped performance pattern concerning where the answer is located within the input. This bias raises concerns, particularly in summarization where crucial content may be dispersed throughout the source document(s). Besides, in summarization, mapping facts from the source to the summary is not trivial as salient content is usually re-phrased. In this paper, we conduct the first comprehensive study on context utilization and position bias in summarization. Our analysis encompasses 6 LLMs, 10 datasets, and 5 evaluation metrics. We introduce a new evaluation benchmark called MiddleSum on the which we benchmark two alternative inference methods to alleviate position bias: hierarchical summarization and incremental summarization [^1] Our code and data can be found here: https://github.com/ntunlp/MiddleSum..
@@ 1 Introduction
Large language models (LLMs) have drastically transformed the landscape of NLP recently Brown et al. (2020). With instruction tuning Ouyang et al. (2022); Chung et al. (2022), LLMs made a major leap forward in conditional (prompted) content generation, and can generate satisfying outputs without the need to fine-tune on a specific task. In abstractive summarization specifically, this approach has arguably opened a new paradigm: summaries generated by LLMs are highly fluent, grammatical and relevant (Goyal et al., 2022). Despite noticeably lower scores on automatic metrics such as ROUGE (Lin, 2004) or BERTScore (Zhang et al., 2019), summaries generated by LLMs are largely preferred by humans over summaries from state-of-the-art fine-tuned models like BRIO (Liu et al., 2022b, 2023c). In fact, on XSum, GPT-3.5 summaries are even on par with re-annotated human-written summaries, and much better than the dataset’s original ground-truth, according to human evaluators (Zhang et al., 2023b). LLMs also show promising capability in evaluating summaries generated by other systems, including LLMs (Fu et al., 2023; Luo et al., 2023; Shen et al., 2023a).
Despite this success, a few major technological bottlenecks remain with LLMs, including the maximum length of their context window. The standard context window length for open-source LLMs is 2k tokens (Brown et al., 2020; Scao et al., 2022; Penedo et al., 2023; Touvron et al., 2023a), which drastically limits their usefulness for long-input summarization (Shaham et al., 2022). Several techniques were proposed to extend the context window, including ALiBi (Press et al., 2021), LeX (Sun et al., 2022), position interpolation (Chen et al., 2023) and YaRN (Peng et al., 2023). While some of them claim up to 100k+ tokens processing capacity (Peng et al., 2023), it remains unclear how much such methods help on long-context summarization.
@@
In this work, we investigate in depth how LLMs use their context window in abstractive summarization. Unlike in question-answering, mapping facts in the output to a specific snippet in the source is not straightforward in abstractive summarization, due to the high-level of re-phrasing and compression. We conduct a large-scale study with 6 LLMs, 10 datasets covering many aspects of summarization, and 5 highly diverse automatic metrics. Our contributions are threefold:
@@ 2 Experimental Setup
We experiment with 6 popular and high-performing LLMs:
  • Flan-UL2 is a 20B parameters encoder-decoder model pre-trained on 1T tokens. It is based on the UL2 20B model (Tay et al., 2022), with the addition of Flan-T5 (Chung et al., 2022) instruction fine-tuning. The context window is 2k tokens.
  • Llama-2 (Touvron et al., 2023b) is a recently introduced powerful decoder-only model pre-trained on 2T tokens, ranging from 7B to 70B parameters, and with a 4k tokens context window. We use the 7B and 13B models.
  • Xgen-7B (Nijkamp et al., 2023) is a 7B decoder-only model pre-trained on up to 1.5T tokens. It supports an 8k tokens context window.
@@
  • GPT-3.5 (gpt-3.5-turbo-0125), which is known to be good at generating summaries and has a 16k tokens context window. [^2] To reduce cost, we subsample 300 data points when using GPT-3.5. Our total API cost is inferior to 300 USD.
@@
We analyze the open-source models through HuggingFace transformers library (Wolf et al., 2020), and use the OpenAI API for GPT-3.5. We use the instruction-tuned (or chat) checkpoints for Llama-2, Xgen-7B and Mistral-7B. Note that popular instruction-tuning datasets such as Flan (Wei et al., 2021) include some of the datasets we study: CNN/DM, XSum, SAMSum and Multi-News. To run inference, we use the following prompt: Read the following text and summarize it: [text]. Summarize the above text in [n] sentences. Summary: where n is set to an average number of target sentences per dataset (see Appendix A). We infer all models in bfloat16 and sample summaries with top-k sampling (Fan et al., 2018) using $k=50$ and temperature $T=0.3$.
@@
Summarization evaluation is especially challenging in the LLM era, as most automatic metrics poorly correlate with human preferences (Goyal et al., 2022; Liu et al., 2023b). To get a broad picture of performance, we evaluate with metrics as diverse as possible. First, we consider reference-based metrics: ROUGE-2 (Lin, 2004), which measures bigram overlap, BERTScore (Zhang et al., 2019), which measures semantic similarity with BERT (Devlin et al., 2019) embeddings, and A3CU (Liu et al., 2023c), which extracts facts in the form of Atomic Content Units (ACUs) Liu et al. (2023b), and checks the presence of ACUs between prediction and reference. As reference-free metrics, we include SummaC (Laban et al., 2022), a leading factual consistency evaluation metric relying on entailment scores between pairs of source and summary sentences. We also leverage GPT-3.5 again (still gpt-3.5-turbo-0125), this time as a summarization evaluator, which is proven to be a strong natural language generation evaluator (Wang et al., 2023a; Shen et al., 2023a; Jain et al., 2023). We prompt the model with the source and generated summary (which fits in GPT-3.5’s 16k context window) and ask to output a score on a likert scale from 1 to 5. We refer to Appendix B for the full prompt template. [^3] We also subsample 300 data points when using GPT-3.5 as an evaluator, to reduce API cost.
We report the performance of LLMs on all 10 datasets, alongside a comparison to SOTA, in Appendix C. FLan-UL2 dominates on standard-length datasets, but GPT-3.5 has the upper hand on the long-input ones. Performance itself is not our focus in this paper, but rather which position-related factors influence it. We discard Flan-UL2 on long-input datasets due to very poor performance.
@@ 3.1 RQ1: Where in the source do LLMs take their information from?
We first investigate summaries generated by LLMs, and map them to specific parts of the input. Unlike in question-answering or extractive summarization, mapping salient information from a summary to the source is not trivial in abstractive summarization.
@@
As seen in Fig. 1, all summarization datasets except XSum, and Reddit-TIFU show some lead bias: salient bigrams from the reference (orange curves) are more likely to be found at the beginning of the source. However, LLMs show a significantly stronger lead bias on all datasets: bigrams from LLMs summaries are much more likely to be found in the first 20% words of the source. It is especially striking on XSum (except for Flan-UL2), Reddit-TIFU, Arxiv, PubMed and GovReport. On XSum, Flan-UL2 closely matches the reference distribution, which we attribute to its better instruction tuning. Results in Appendix D confirm that bigram distribution for LLMs and references are statistically different (p-value of Kolmogorov-Smirnov test (Massey Jr, 1951) inferior to 0.001) in all but 4 out of the 55 (dataset, LLM) setups: Flan-UL2 on XSum and SAMSum, Llama-2-7B on SAMSum and GPT-3.5 on SAMSum. We conclude that LLMs focus on contents at the beginning of the source document(s).
@@ 3.2 RQ2: Where do LLMs look at within their context window?
In the previous experiment’s design, LLMs may not see the entire source in long-input summarization datasets, due to their limited context window, which is shorter than the source on the long-input datasets. We now focus on input information accessible to LLMs, and only consider salient information if it falls within the context window. Besides, since the same bigram may occur multiple times throughout the source, we adjust the methodology for saliency estimation. We align sentences in generated summaries to sentences in the context, following the procedure described in (Zhou et al., 2018) and also used in (Adams et al., 2023). Specifically, we greedily select source sentences (among the ones fitting in context window) until the ROUGE-1 F1 score between the set of selected source sentences and the summary stops increasing. The resulting set of source sentences forms a proxy of the visible salient input information being rephrased by the model when summarizing. We split each truncated source document into 10 bins of the same number of sentences, and map each aligned source sentence to its bin. Note that bins are not directly comparable across models, as context length varies across models.
As we can see in Fig. 2, sentences from the first 10% or last 10% of the input context are much more represented than others. A clear U-shape emerges on PubMed and SummScreenFD for all LLMs. This is intriguing knowing that the LLMs have different context window lengths, and the last 10% of each context window may contain content of varying saliency. In other words, LLMs seem to be mostly re-phrasing information from the beginning or the end of their context window.
Results in Appendix E confirm an even stronger position bias with base models.
@@ 3.3 RQ3: Does LLMs performance depend on the position of salient information?
We examine performance changes with regard to this salient position. To do so, we compute the Spearman correlation coefficient between salient position and each evaluation metric in Table 2. A high absolute Spearman value means that summary quality (as measured by this metric) can change (and deteriorate) with the position of important information within the context.
There are several takeaway findings from this Table. First, we notice that on standard-length datasets, reference-based evaluation metrics are negatively correlated to position of salient information. The correlation is only moderate, yet remarkably consistent across datasets (except Reddit-TIFU) and models. This is surprising, since such datasets fit entirely in context and are not affected by truncation. In contrast, reference-free metrics show either no significant or positive correlation to information position. For long-input datasets, the negative trend for reference-based metrics is confirmed. On these lengthy datasets, SummaC and GPT-3.5 tend to switch from positive to negative correlation, especially for Xgen-7B and Mistral-7B. We highlight that since GPT-3.5 itself is affected by the middle-curse from Liu et al. (2023a), it may not accurately evaluate summarization when salient content lays in the middle of the context. In light of these results, we conservatively conclude that LLMs’ summarization performance is sensitive to the position of salient information in the context window.
@@ 4.1 How is information in the middle treated?
In the first experiment, we vary the position of salient information throughout the input. We keep a single document (the abstract of the query paper on Multi-XScience, and the document with the highest BERTScore with the reference on Multi-News), and place it at position $j$ for $j\in\{1,\ldots,k\}$, using $k-1$ documents from a random data point for the other slots. The single relevant document is accompanied by a [RELEVANT] header, while the other documents have an [IRRELEVANT] header, and we prompt the LLM to only summarize the relevant document. For reference-free evaluation metrics, we use the single relevant document as source. We also include a random baseline of shuffled inputs and model predictions. In Fig. 3, we see a noticeable drop in performance for all metrics when the salient document is not in the first or final position. Flan-UL2 seems to focus on the end of the context, Xgen-7B and Mistral-7B on the beginning, and Llama-2 models and GPT-3.5 on both. Performance can fall quite below random range, especially for reference-free metrics, confirming the worrying trend from Liu et al. (2023a).
A more fine-grained analysis with GPT-3.5 in Fig. 4 evaluating specific attributes (following the method in Adams et al. (2023), see Appendix B) reveals more details. Coherence and quality remain high and stable. In other words, the text outputs from the LLMs are always of good quality. But for informativeness and attributability, the U-shape appears again: it shows that the LLMs (even the powerful GPT-3.5) are struggling to generate content specifically sticking to the document inserted in the middle.
Dataset Model Input documents R-2 BS A3CU SummaC GPT-3.5 %
Multi-X Llama-2-7B All 7 4.64 82.83 5.88 54.56 4.17 100.00
First + last 4.62 82.82 5.78 47.00 4.50 98.36
First + 5 random + last 4.43 82.64 5.37 43.25 4.01 92.40
@@
First + last 4.73 82.86 5.76 43.74 4.53 98.46
First + 5 random + last 4.61 82.80 5.72 46.42 4.31 98.07
@@
First + last 5.01 82.73 5.86 49.08 4.45 99.83
First + 5 random + last 3.89 82.16 5.03 55.29 3.01 88.93
@@
First + last 5.12 82.67 6.15 60.91 4.76 99.80
First + 5 random + last 4.45 82.40 5.35 58.19 4.06 90.59
@@
First + last 4.75 83.04 6.05 39.81 4.69 96.42
First + 5 random + last 4.37 82.90 5.67 43.26 4.63 95.63
Multi-N Llama-2-7B All 5 10.76 85.04 19.06 60.09 4.00 100.00
First + last 9.50 84.43 15.88 54.52 3.80 91.32
First + 3 random + last 7.57 83.36 12.39 50.35 2.94 78.13
@@
First + last 9.55 84.58 16.99 49.84 3.73 93.93
First + 3 random + last 8.27 83.79 14.84 50.09 3.18 86.14
@@
First + last 7.82 83.27 14.18 51.59 3.60 92.68
First + 3 random + last 6.30 81.85 11.66 49.02 2.66 79.51
[…]
First + last 9.11 83.69 14.99 67.14 3.51 100.37
First + 3 random + last 6.59 81.66 12.50 52.14 2.45 80.17
[…]
First + last 8.94 84.52 15.85 45.94 4.01 92.76
First + 3 random + last 8.53 84.31 15.33 44.67 3.80 89.81
Table 3: Performance in multi-document summarization on Multi-XScience (7 documents) and Multi-News (5 documents) when infilling the middle of the context window with random documents. R-2 is ROUGE-2, BS refers to BERTScore. % is the mean relative performance compared to the baseline with all documents.
@@
In the second experiment, we take the opposite approach, and put salient information at the beginning and the end, while the middle of the prompt is filled with noise. We keep the first and last documents, and fill the $k-2$ middle ones with random documents. We also run a baseline just using the first and last documents as input, expected to be close to the result with random documents in between. As displayed in Table 3, filling with random noise between the first and last document (which amounts to a prompt mostly irrelevant to the reference) leads to a moderate drop in performance. For instance, on Multi-XScience, with 5 random documents between the first and last, Llama-2-13B maintains 98% of its performance, and reaches a GPT-3.5 score of 4.31 as compared to 4.35 when using all 7 documents.